OpenStudy (anonymous):

Find 3 quadratic equations that go through (0,0) and (6,0). I have found two. I cannot find a third. 1. f(x)=x(x-6) 2. f(x)= -x(x-6)

OpenStudy (math&ing001):

This one works too f(x)=x^2(x-6)

OpenStudy (anonymous):

Thank you, but it needs to be a parabola. I appreciate your input though!

OpenStudy (amistre64):

given 2 points, there are an infinite number of quadratics that wil fit

OpenStudy (amistre64):

define a vertex; (a,b) to build the vertex format with y = k(x-a)^2+b now for any given points: (p,q) and (s,t) we can setup a system of equations q = k(p-a)^2+b t = k(s-a)^2+b

OpenStudy (amistre64):

or do a matrix run

OpenStudy (jdoe0001):

notice that your x-intercepts, thhat is, your endpoints are 

OpenStudy (amistre64):

these are roots arent they :/

OpenStudy (amistre64):

given 2 roots: a,b y = k(a-x)(b-x) satisfies all quadratics with roots a,b for any real k

OpenStudy (jdoe0001):

so use the parabola "vertex form" and pick a point, as @amistre64 said, between those 2 given points for the vertex once you get those coordinates for the vertex between those 2 points, plug in your ( 6, 0) values and solve for "a" \(\bf\large y = \color{red}{a}(x-\textit{x-coord here})^2+\textit{y-coord here}\)

OpenStudy (jdoe0001):

say for example, let's pick... say between 0 ... 6 for "x" say.... (3, -5) down below \(\bf y = \color{red}{a}(x-3)^2-5\qquad x = 6\qquad y = 0\qquad (6,0)\qquad thus\\ \quad \\ y = \color{red}{a}(x-3)^2-5\implies 0 = \color{red}{a}(6-3)^2-5\)

OpenStudy (jdoe0001):

as and @amistre64 said..... over teh x=3 line, you can pick an infinite values for "y" to get infinite amount of vertices and thus infinite parabola functions

OpenStudy (math&ing001):

Bottom line f(x)=2x(x-6) will do

OpenStudy (amistre64):

im not sure i like that bottom line ... its just to flat